# ShieldAgent: Shielding Agents via Verifiable Safety Policy Reasoning

### v1

@@ 3.3 ShieldAgent Framework
Tool Library. To support these operations, ShieldAgent is equipped with powerful tools, including moderation APIs for various modalities (e.g., image, video, audio) and formal verification tools (e.g., Stormpy). To enhance guardrail accuracy, we fine-tuned two specialized guardrail models based on InternVL2-2B (Chen et al., 2024b) for enumeration-based search and binary-check operations.
@@
Built on the MCP framework (Anthropic, 2024), ShieldAgent collectively integrates these modules to handle diverse shielding scenarios while allowing users to customize new tools to extend the guardrail capabilities.
@@ 4 ShieldAgent-Bench Dataset
Figure 2: Pipeline for curating ShieldAgent-Bench. We adopt the AWM web agent (Wang et al., 2024) and collect safe trajectories by executing instructions with full policy compliance. For risky trajectories, we attack the agent with two SOTA agent-based and environment-based algorithms and produce unsafe trajectories across seven risk categories.
@@
Safety-related Instructions. We selectively reuse the instruction templates from WebArena (Zhou et al., 2023) and ST-WebAgentBench (Levy et al., 2024) across six environments (i.e., Shopping, CMS, Reddit, GitLab, Maps, SuiteCRM), and curate instructions that yield potential safety risks by augmenting the templates with safety-critical information (e.g. API token). Finally, we obtain 960 high-quality safety-related instructions. Specifically, each sample in our dataset consists of $(I_{s},\zeta_{s},\zeta_{u}^{a},\zeta_{u}^{e})$, where $I_{s}$ is the instruction, $\zeta_{s}$ is the safe trajectory, and $\zeta_{u}^{a}$, $\zeta_{u}^{e}$ are unsafe trajectories induced by two types of attacks, respectively. Each $\zeta$ includes the complete interactions between the agent and the environment at each step, including: (1) all conversations, (2) visual screenshots, (3) HTML accessibility trees.
Policy-Targeted Agent Attacks. We consider two types of adversarial perturbations against agents, each instanced by a practical attack algorithm: (1) Agent-based: we adopt AgentPoison (Chen et al., 2024c), which injects adversarial demonstrations in the agent’s memory or knowledge base to manipulate its decision-making; (2) Environment-based: we adopt AdvWeb (Xu et al., 2024), which stealthily manipulates the environment elements to mislead the agent. Specifically, we adapt both algorithms to attack a SOTA web agent, AWM (Wang et al., 2024) to violate at least one extracted safety policy per instruction, ensuring policy-centered safety violation for tractable guardrail evaluation.
Comprehensive Risk Categories. We carefully investigate the extracted policies, risky trajectories induced by our attack, and concurrent studies on agents’ risky behaviors (Levy et al., 2024), and categorize the unsafe trajectories into seven risk categories: (1) access restriction, (2) content restriction, (3) hallucination, (4) instruction adherence, (5) operational restriction, (6) typical error patterns, and (7) long-term risks. Please refer to Appendix F for more details.
@@ 5.1 Setup
Datasets. We evaluate ShieldAgent against guardrail baselines on our ShieldAgent-Bench dataset and three existing benchmarks: (1) ST-WebAgentBench (Levy et al., 2024), which includes 234 safety-related web agent tasks with simple safety constraints; (2) VWA-Adv (Wu et al., 2025), consisting of 200 realistic adversarial tasks in the VisualWebArena (Koh et al., 2024); and (3) AgentHarm (Andriushchenko et al., ), comprising 110 malicious tasks designed for general agents. Notably, to properly evaluate agent guardrails, each sample must include an instruction, agent trajectory, enforced policy, and ground-truth label as protocols—all of which are available in ShieldAgent-Bench. However, existing benchmarks only provide task instructions (see Table 1). To address this, we augment them by collecting corresponding policies and both safe and unsafe trajectories using various algorithms. See Appendix F for details on the curation pipeline and dataset statistics.
Baselines. We consider three representative baselines: (1) Direct prompt: We provide GPT-4o with the complete policy and directly prompt it to produce an overall safety label and any violated rules. (2) Rule traverse: We traverse each rule and prompt GPT-4o to identify potential violation. We flag the trajectory as unsafe once a rule is flagged as violated. (3) GuardAgent (Xiang et al., 2024): We follow their pipeline and set the guard request to identify any policy violations in the agent trajectory. To ensure a fair comparison, we provide all methods with the same safety policy as input and collect the following outputs for evaluation: (i) A binary flag (safe or unsafe); (ii) A list of violated rules, if any.
@@ 5.2 Results
Online Guardrail. We further evaluate ShieldAgent’s performance in providing online guardrails for web agents. Specifically, we use the AWM agent as the task agent and integrate each guardrail method as a post-verification module that copilots with the agent. These guardrails verify the agent’s actions step-by-step and provide interactive feedback to help it adjust behavior for better policy compliance. Notably, this evaluation setting comprehensively captures key dimensions such as guardrail accuracy, fine-grained policy grounding, and explanation clarity, which are all critical components for effectively guiding the task agent’s behavior toward better safety compliance. As shown in Table 4, ShieldAgent also outperforms all baselines in this online setting, achieving the highest policy compliance rate. These results highlight ShieldAgent’s effectiveness as System 2 (Li et al., 2025) to seamlessly integrate with task agents to enhance their safety across diverse environments.
@@ Acknowledgment
This work is partially supported by the National Science Foundation under grant No. 1910100, No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA TIAMAT No. 80321, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, Alfred P. Sloan Fellowship, the research grant from eBay, AI Safety Fund, Virtue AI, and Schmidt Science.
@@ A.1 Notations
Let $\mathcal{X}$ denote the environment, and let $\pi_{\text{agent}}$ be the action policy of an agent we aim to shield. At each step $i$, the agent receives an observation $o_{i}\in\mathcal{X}$ and maps it to a partial state $s_{i}=f(o_{i})$ via a state-space mapping function $f$. Specifically for web agents, $f$ extracts accessibility trees (AX-trees) from the webpage’s HTML and visual screenshots, condensing key information from lengthy observations (Zhou et al., 2023). Then, the agent generates an action $a_{i}$ by sampling from policy $a_{i}\sim\pi_{\text{agent}}(s_{i})$ and progressively interacts with the environment $\mathcal{X}$.
@@ A.2 Solution Space
Since predicates can sometimes be inaccurately assigned, $\phi_{p}$ can serve as knowledge in ASPM to enhance the robustness of our shield (Kang & Li, 2024). With these rules, ShieldAgent can effectively reason in the solution space to shield the agent action with high accuracy and robustness.
@@ B.1 ST-WebAgentBench
Table 5: Comparison of guardrail performance across three risk categories in ST-WebAgentBench (Levy et al., 2024). Specifically, we report the averaged accuracy (ACC) and false positive rate (FPR) for each evaluation category, along with overall averages. The best performance is in bold.
@@ B.2 VWA-Adv
Specifically, VWA-Adv (Wu et al., 2025) attacks web agents by perturbing either the text instruction by adding a suffix or the image input by adding a bounded noise. Specifically, VWA-Adv constructs 200 diverse risky instructions based on the three environments from VisualWebArena (Koh et al., 2024). The environments are detailed as follows:
@@
Shopping. This environment builds on the e-commerce site from WebArena (Zhou et al., 2023), where successful navigation requires both textual and visual comprehension of product images, reflecting typical online shopping tasks.
@@
Table 6: Guardrail performance comparison on VWA-Adv (Wu et al., 2025) across three environments in VisualWebArena (Koh et al., 2024), i.e., Classifieds, Reddit, Shopping, under two perturbation sources, i.e., text-based and image-based. We report accuracy (ACC) and false positive rate (FPR) for each environment. The best performance is in bold.
@@ B.3 AgentHarm
Table 7: Guardrail performance comparison on AgentHarm (Andriushchenko et al., ) across 11 harm categories. The reported numbers have been slightly perturbed (by ±$\pm 0.1$ in an alternating pattern) and the overall is recalculated as the arithmetic mean of the 11 categories. The best performance is in bold.
@@ C.3 Linear Temporal Logic (LTL) Rules
Temporal logic represents propositional and first-order logical reasoning with respect to time. Linear temporal logic over finite traces ($\text{LTL}_{f}$(Zhu et al., 2017) is a form of temporal logic that deals with finite sequences, i.e., finite-length trajectories.
@@ C.4 ASPM Structure Optimization
0:  Predicate set $\mathcal{P}=\{\mathcal{P}_{a},\mathcal{P}_{s}\}$; Rule set =$\mathcal{R}=\{\mathcal{R}_{a},\mathcal{R}_{p}\}$; Embedding model $\mathcal{E}$; Clustering algorithm $\mathcal{C}$; Refinement budget $N_{\mathrm{b}}$; Max iterations $M_{\mathrm{it}}$; Surrogate LLM; Graph $G=(\mathcal{P},E)$ with initial edge weights $E$.
1:  Initialize vagueness score for each predicate $\mathcal{V}_{p},p\in\mathcal{P}$triangleright Calculate via Eq. (3)
2:  $\mathcal{V}_{r}=\max\{\mathcal{V}_{p_{1}},\dots,\mathcal{V}_{p_{|\mathcal{P}_{r}|}}\},\mathcal{P}_{r}\subseteq\mathcal{P}$triangleright Compute vagueness score for each rule
@@
4:  $n\leftarrow 0$triangleright Count how many refinements have been done
5:  for $m=1\ \text{to}\ M_{\mathrm{it}}$ do
@@
8:        (_$(\_,r)\leftarrow\mathrm{HeapPop}(\mathcal{U})$triangleright Pop the most vague rule
@@
10:           $r_{\mathrm{new}}\leftarrow\mathrm{LLM\_refine}\bigl{(}r,\ \mathcal{P}_{r}\bigr{)}$triangleright Refine rule $r$ to be verifiable; update its predicates if needed
@@
20:     $\mathcal{K}\leftarrow\mathcal{C}(G)$triangleright Cluster predicates in $G$ to prune redundancy
@@
22:        $p_{\mathrm{merged}}\leftarrow\mathrm{LLM\_merge}\bigl{(}C,\ \mathcal{R}\bigr{)}$triangleright Merge similar predicates/rules in $C$ if beneficial
@@ C.5 Training ASPM
1:  $A\in\{0,1\}^{|\mathcal{P}_{s}|\times|\mathcal{P}_{s}|}\leftarrow\mathbf{0}$triangleright Initialize adjacency matrix
2:  $A_{ij}\leftarrow 1\ \mathrm{if}\ (p^{i}_{s},p^{j}_{s})\text{ co-occur in any rule}\ \mathrm{OR}\ \mathrm{cosSim}\bigl{(}\mathrm{emb}(p^{i}_{s}),\,\mathrm{emb}(p^{j}_{s})\bigr{)}\!\geq\!\theta;\ \text{else }0.$triangleright Build adjacency matrix
3:  $\mathrm{labels}\leftarrow\textsc{SpectralClustering}(A,\,k)$triangleright Cluster the state predicates into $k$ groups
@@
5:     $C^{\ell}_{p}\leftarrow\{p_{s}\mid\mathrm{labels}[p_{s}]=\ell\}$triangleright Form predicate clusters $\mathcal{C}_{p}$
@@
9:        $\mathcal{C}^{\ell}_{p}\leftarrow\mathcal{C}^{\ell}_{p}\cup\mathcal{C}^{m}_{p}\,\,\text{s.t.}\,\,p_{s}^{i}\in\mathcal{C}^{\ell}_{p},p_{s}^{j}\in\mathcal{C}^{m}_{p}$triangleright If two co-occurring predicates appear in different clusters, merge them
@@
13:     $C^{\ell}_{r}\leftarrow\{r_{s}\mid p_{s}\in C^{\ell}_{p}\}$triangleright Group rules which share state predicates in the same cluster
@@
15:  $\mathcal{G}_{\text{ASPM}}\leftarrow\varnothing$triangleright Initialize ASPM as an empty dictionary with actions as keys
@@
20:              $\mathcal{G}_{\text{ASPM}}[p_{a}]=\mathcal{G}_{\text{ASPM}}[p_{a}]\cup C^{\ell}_{r}$triangleright Associate action circuits with any relevant rule clusters
@@
31:        for $i=1$ to $N$ do
32:           Compute $P_{\theta}(\mu_{p_{a}=1}^{(i)})$ and $P_{\theta}(\mu_{p_{a}=0}^{(i)})$triangleright Run probabilistic inference to obtain corresponding safety probabilities via Eq. (4)
33:           Compute loss $\mathcal{L}(\theta)$triangleright Calculate loss w.r.t. the groundtruth labels via Eq. (6)

### v2

@@ 3.3 ShieldAgent Framework
Tool Library. To support these operations, ShieldAgent is equipped with powerful tools, including moderation APIs for various modalities (e.g., image, video, audio) and formal verification tools (e.g., Stormpy). To enhance guardrail accuracy, we fine-tuned two specialized guardrail models based on InternVL2-2B (chen2024far) for enumeration-based search and binary-check operations.
@@
Built on the MCP framework (mcp2025), ShieldAgent collectively integrates these modules to handle diverse shielding scenarios while allowing users to customize new tools to extend the guardrail capabilities.
@@ 4 ShieldAgent-Bench Dataset
Figure 2: Pipeline for curating ShieldAgent-Bench. We adopt the AWM web agent (wang2024agent) and collect safe trajectories by executing instructions with full policy compliance. For risky trajectories, we attack the agent with two SOTA agent-based and environment-based algorithms and produce unsafe trajectories across seven risk categories.
@@
Safety-related Instructions. We selectively reuse the instruction templates from WebArena (zhou2023webarena) and ST-WebAgentBench (levy2024st) across six environments (i.e., Shopping, CMS, Reddit, GitLab, Maps, SuiteCRM), and curate instructions that yield potential safety risks by augmenting the templates with safety-critical information (e.g. API token). Finally, we obtain 960 high-quality safety-related instructions. Specifically, each sample in our dataset consists of $(I_{s},\zeta_{s},\zeta_{u}^{a},\zeta_{u}^{e})$, where $I_{s}$ is the instruction, $\zeta_{s}$ is the safe trajectory, and $\zeta_{u}^{a}$, $\zeta_{u}^{e}$ are unsafe trajectories induced by two types of attacks, respectively. Each $\zeta$ includes the complete interactions between the agent and the environment at each step, including: (1) all conversations, (2) visual screenshots, (3) HTML accessibility trees.
Policy-Targeted Agent Attacks. We consider two types of adversarial perturbations against agents, each instanced by a practical attack algorithm: (1) Agent-based: we adopt AgentPoison (chenagentpoison), which injects adversarial demonstrations in the agent’s memory or knowledge base to manipulate its decision-making; (2) Environment-based: we adopt AdvWeb (xu2024advweb), which stealthily manipulates the environment elements to mislead the agent. Specifically, we adapt both algorithms to attack a SOTA web agent, AWM (wang2024agent) to violate at least one extracted safety policy per instruction, ensuring policy-centered safety violation for tractable guardrail evaluation.
Comprehensive Risk Categories. We carefully investigate the extracted policies, risky trajectories induced by our attack, and concurrent studies on agents’ risky behaviors (levy2024st), and categorize the unsafe trajectories into seven risk categories: (1) access restriction, (2) content restriction, (3) hallucination, (4) instruction adherence, (5) operational restriction, (6) typical error patterns, and (7) long-term risks. Please refer to Appendix F for more details.
@@ 5.1 Setup
Datasets. We evaluate ShieldAgent against guardrail baselines on our ShieldAgent-Bench dataset and three existing benchmarks: (1) ST-WebAgentBench (levy2024st), which includes 234 safety-related web agent tasks with simple safety constraints; (2) VWA-Adv (wudissecting), consisting of 200 realistic adversarial tasks in the VisualWebArena (koh2024visualwebarena); and (3) AgentHarm (andriushchenkoagentharm), comprising 110 malicious tasks designed for general agents. Notably, to properly evaluate agent guardrails, each sample must include an instruction, agent trajectory, enforced policy, and ground-truth label as protocols—all of which are available in ShieldAgent-Bench. However, existing benchmarks only provide task instructions (see Table 1). To address this, we augment them by collecting corresponding policies and both safe and unsafe trajectories using various algorithms. See Appendix F for details on the curation pipeline and dataset statistics.
Baselines. We consider three representative baselines: (1) Direct prompt: We provide GPT-4o with the complete policy and directly prompt it to produce an overall safety label and any violated rules. (2) Rule traverse: We traverse each rule and prompt GPT-4o to identify potential violation. We flag the trajectory as unsafe once a rule is flagged as violated. (3) GuardAgent (xiang2024guardagent): We follow their pipeline and set the guard request to identify any policy violations in the agent trajectory. To ensure a fair comparison, we provide all methods with the same safety policy as input and collect the following outputs for evaluation: (i) A binary flag (safe or unsafe); (ii) A list of violated rules, if any.
@@ 5.2 Results
Online Guardrail. We further evaluate ShieldAgent’s performance in providing online guardrails for web agents. Specifically, we use the AWM agent as the task agent and integrate each guardrail method as a post-verification module that copilots with the agent. These guardrails verify the agent’s actions step-by-step and provide interactive feedback to help it adjust behavior for better policy compliance. Notably, this evaluation setting comprehensively captures key dimensions such as guardrail accuracy, fine-grained policy grounding, and explanation clarity, which are all critical components for effectively guiding the task agent’s behavior toward better safety compliance. As shown in Table 4, ShieldAgent also outperforms all baselines in this online setting, achieving the highest policy compliance rate. These results highlight ShieldAgent’s effectiveness as System 2 (li2025system) to seamlessly integrate with task agents to enhance their safety across diverse environments.
@@ Acknowledgment
We thank Meng Ding for the constructive suggestions and help with the paper writing. This work is partially supported by the National Science Foundation under grant No. 1910100, No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA TIAMAT No. 80321, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, Alfred P. Sloan Fellowship, the research grant from eBay, AI Safety Fund, Virtue AI, and Schmidt Science.
@@ A.1 Notations
Let $\mathcal{X}$ denote the environment, and let $\pi_{\text{agent}}$ be the action policy of an agent we aim to shield. At each step $i$, the agent receives an observation $o_{i}\in\mathcal{X}$ and maps it to a partial state $s_{i}=f(o_{i})$ via a state-space mapping function $f$. Specifically for web agents, $f$ extracts accessibility trees (AX-trees) from the webpage’s HTML and visual screenshots, condensing key information from lengthy observations (zhou2023webarena). Then, the agent generates an action $a_{i}$ by sampling from policy $a_{i}\sim\pi_{\text{agent}}(s_{i})$ and progressively interacts with the environment $\mathcal{X}$.
@@ A.2 Solution Space
Since predicates can sometimes be inaccurately assigned, $\phi_{p}$ can serve as knowledge in ASPM to enhance the robustness of our shield (kang2024r). With these rules, ShieldAgent can effectively reason in the solution space to shield the agent action with high accuracy and robustness.
@@ B.1 ST-WebAgentBench
Table 5: Comparison of guardrail performance across three risk categories in ST-WebAgentBench (levy2024st). Specifically, we report the averaged accuracy (ACC) and false positive rate (FPR) for each evaluation category, along with overall averages. The best performance is in bold.
@@ B.2 VWA-Adv
Specifically, VWA-Adv (wudissecting) attacks web agents by perturbing either the text instruction by adding a suffix or the image input by adding a bounded noise. Specifically, VWA-Adv constructs 200 diverse risky instructions based on the three environments from VisualWebArena (koh2024visualwebarena). The environments are detailed as follows:
@@
Shopping. This environment builds on the e-commerce site from WebArena (zhou2023webarena), where successful navigation requires both textual and visual comprehension of product images, reflecting typical online shopping tasks.
@@
Table 6: Guardrail performance comparison on VWA-Adv across three environments in VisualWebArena, i.e., Classifieds, Reddit, Shopping, under two perturbation sources, i.e., text-based and image-based. We report accuracy (ACC) and false positive rate (FPR) for each environment. The best performance is in bold.
@@ B.3 AgentHarm
Table 7: Guardrail performance comparison on AgentHarm across 11 harm categories. The best performance is in bold.
@@ C.3 Linear Temporal Logic (LTL) Rules
Temporal logic represents propositional and first-order logical reasoning with respect to time. Linear temporal logic over finite traces ($\text{LTL}_{f}$(zhu2017symbolic) is a form of temporal logic that deals with finite sequences, i.e., finite-length trajectories.
@@ C.4 ASPM Structure Optimization
0: Predicate set $\mathcal{P}=\{\mathcal{P}_{a},\mathcal{P}_{s}\}$; Rule set $\mathcal{R}=\{\mathcal{R}_{a},\mathcal{R}_{p}\}$; Embedding model $\mathcal{E}$; Clustering algorithm $\mathcal{C}$; Refinement budget $N_{\mathrm{b}}$; Max iterations $M_{\mathrm{it}}$; Surrogate LLM; Graph $G=(\mathcal{P},E)$ with initial edge weights $E$.
1: Initialize vagueness score for each predicate $\mathcal{V}_{p},p\in\mathcal{P}$ $\triangleright$ Calculate via Eq. (3)
[…]
@@
[…]
@@
8:   $(\_,r)\leftarrow\mathrm{HeapPop}(\mathcal{U})$ $\triangleright$ Pop the most vague rule
@@
10:    $r_{\mathrm{new}}\leftarrow\mathrm{LLM\_refine}\bigl(r,\ \mathcal{P}_{r}\bigr)$ $\triangleright$ Refine rule $r$ to be verifiable; update its predicates if needed
@@
20:  $\mathcal{K}\leftarrow\mathcal{C}(G)$ $\triangleright$ Cluster predicates in $G$ to prune redundancy
@@
22:   $p_{\mathrm{merged}}\leftarrow\mathrm{LLM\_merge}\bigl(C,\ \mathcal{R}\bigr)$ $\triangleright$ Merge similar predicates/rules in $C$ if beneficial
@@
Table 8: Statistics of ASPM before and after policy model structure optimization across each environment. Specifically, we demonstrate the number of predicates, the number of rules, and the average vagueness score of each rule. The maximum number of iterations is set to 10 across all environments.
Environment Before Optimization After Optimization
# Predicates # Rules Avg. Vagueness # Predicates # Rules Avg. Vagueness
Shopping 920 562 0.71 461 240 0.38
CMS 590 326 0.69 225 120 0.34
Reddit 1150 730 0.77 490 178 0.49
GitLab 1079 600 0.62 363 198 0.51
Maps 430 202 0.64 210 104 0.25
SuiteCRM 859 492 0.66 390 240 0.32
@@ C.5 Training ASPM
[…]
@@
5:  $C^{\ell}_{p}\leftarrow\{p_{s}\mid\mathrm{labels}[p_{s}]=\ell\}$ $\triangleright$ Form predicate clusters $\mathcal{C}_{p}$
@@
9:   $\mathcal{C}^{\ell}_{p}\leftarrow\mathcal{C}^{\ell}_{p}\cup\mathcal{C}^{m}_{p}\,\,\text{s.t.}\,\,p_{s}^{i}\in\mathcal{C}^{\ell}_{p},p_{s}^{j}\in\mathcal{C}^{m}_{p}$ $\triangleright$ If two co-occurring predicates appear in different clusters, merge them
@@
13:  $C^{\ell}_{r}\leftarrow\{r_{s}\mid p_{s}\in C^{\ell}_{p}\}$ $\triangleright$ Group rules which share state predicates in the same cluster
@@
[…]
@@
20:     $\mathcal{G}_{\text{ASPM}}[p_{a}]=\mathcal{G}_{\text{ASPM}}[p_{a}]\cup C^{\ell}_{r}$ $\triangleright$ Associate action circuits with any relevant rule clusters
@@
31:   for $i=1$ to NN do
32:    Compute $P_{\theta}(\mu_{p_{a}=1}^{(i)})$ and $P_{\theta}(\mu_{p_{a}=0}^{(i)})$ $\triangleright$ Run probabilistic inference to obtain corresponding safety probabilities via Eq. (4)
33:    Compute loss $\mathcal{L}(\theta)$ $\triangleright$ Calculate loss w.r.t. the groundtruth labels via Eq. (6)